# Gaussian Processes enabled model calibration in the context of deep geological disposal

### v1

@@ 5.2 GP-based surrogate and sensitivity analysis
Figure 7: Time-dependent global sensitivity analysis. First and total order Sobol’ indices over time for vertical (left column) and horizontal (right column) convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), presented from top to bottom, respectively.
@@
Figure 8: Aggregated global sensitivity analysis. Normalized first-order (top row) and total-order (bottom row) Sobol’ indices for vertical (left column) and horizontal (right column) convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), from left to right, respectively. The cumulated first-order and total Sobol’ indices are normalized against their respective highest overall values for fair comparison with the other indicators.
@@ 5.3 Model calibration from in-situ monitoring data
Figure 9: Model calibration with monitoring data. Comparison of monitoring data (black squares), GP-based optimal surrogate model prediction (red dots), and corresponding FLAC3D simulation results using the optimal parameter values (blue line) for the vertical (left) and horizontal (right) convergences.
@@ 6 Conclusion and Outlook
We presented an automated workflow that enables repeated online calibration of monitoring data, contributing to digital twinning in the context of deep geological disposals. The considered numerical model represents the mechanical behavior of an emplacement drift, and the calibration from real-world monitoring data has shown excellent results. First, a sensitivity analysis with $16,384$ surrogate model calls was performed within $4.5$ minutes to identify four relevant material parameters. Afterwards, the GP-based surrogate model was calibrated based on convergence measurements, including the convergence of the global optimization in $29$ iterations, $1805$ model evaluations for gradient construction, and a duration of less than $1$ second.
We have identified three key action points for future work. First, the GP calibration and optimization method must be extended to account for higher-dimensional monitoring data, which could include extensometer and permeability measurements. Second, an alternative formulation would be developing a surrogate model that considers the time dependency, which enables forecasting capabilities together with uncertainty propagation. Third, efforts have to be undertaken to account for the multi-physics nature of deep geological disposal. Apart from the purely mechanical model investigated here, a variety of models has been developed in the past that describe, e.g., transport of radio-nuclides through fluid flow and heat generation as a result of disposal of high-level radioactive waste or hydration of a sealing structure. Due to the large difference in time and length scales of the aforementioned models, a monolithic coupling is infeasible. Alternatively, we plan to investigate to which extent GP-based surrogates can be embedded into a multi-physics workflow, possibly accompanied by agent systems.
@@ Appendix A Appendix - Preliminary parameter selection
A preliminary parameter selection was performed in order to identify relevant strain components of TUBSsalt given the considered monitoring data. For this purpose, several damage-associated strain components were deactivated, in particular tertiary creep $\dot{\varepsilon}_{t}$, creep and shear failure $\dot{\varepsilon}_{n}$ and tension failure $\dot{\varepsilon}_{z}$. A comparison between $274$ gridpoint displacements at $40$ time instances is given in Figure 10 below, in which the vertical (right) and horizontal (left) displacements were simulated once with and once without damage strains. From the comparison, it can be observed that no deviation from the diagonal line is visible, which is also confirmed by the $R^{2}$ value close to $1.0$. Thus, the considered test case is dominated by creep mechanisms and softening, and post-failure strains are considered negligible.
@@ Appendix B Appendix - Material parameter
Table 1 contains parameter values and ranges for the constitutive model TUBSsalt. As reference serves a parameter set determined for Gorleben salt in [48]. However, it should be noted that the parameter set has been identified using data from various salt formations of Gorleben, and not only from the homogeneous z2HS2 area. As no parameters for healing are given in [48], $\eta_{v}$ and $m_{v}$ are obtained from [10].
Table 1: TUBSsalt material parameter under consideration of reference values from [48]. The parameters highlighted in red are directly read from experimental data. For the parameters highlighted in blue, a range of values is indicated in order to perform the sensitivity analysis.
Symbol Description Value or value range Unit

### v2

@@ 5.2 GP-based surrogate and sensitivity analysis
Figure 7: Time-dependent global sensitivity analysis. First and total order Sobol’ indices over time for a vertical and b horizontal convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), presented from top to bottom, respectively.
@@
Figure 8: Aggregated global sensitivity analysis. Normalized first-order (a - b) and total-order (c - d) Sobol’ indices for vertical (a - c) and horizontal (b - d) convergences, showing the influence of the primary creep parameters ($\eta_{p}$, $\sigma_{0,eq,p}$, $p_{p}$, $E_{p}$) and secondary creep parameters ($\eta_{s}$, $\sigma_{0,eq,s}$, $p_{s}$), from left to right, respectively. The cumulated first-order and total Sobol’ indices are normalized against their respective highest overall values for fair comparison with the other indicators.
@@ 5.3 Model calibration from in-situ monitoring data
Figure 9: Model calibration with monitoring data. Comparison of monitoring data (black squares), GP-based optimal surrogate model prediction (red dots), and corresponding FLAC3D simulation results using the optimal parameter values (blue line) for a. the vertical and b. horizontal convergences.
@@ 6 Conclusion and Outlook
Deep geological disposal of hazardous materials requires robust numerical models to ensure long-term safety and stability. The calibration of such models with real-world monitoring data is essential for accurately reflecting in-situ conditions and enhancing repository management. This study contributes to advancing digital twinning for deep geological disposals by automating the calibration process using Gaussian Process (GP)-based surrogate models.
The presented workflow combines sensitivity analysis, surrogate modeling, and optimization to enable efficient calibration of a mechanical model representing the behavior of an emplacement drift in rock salt formations located in the northern main drift of Gorleben, Germany. Sensitivity analysis was conducted using time-dependent Sobol’ indices to identify key material parameters, followed by calibration of the sensitive parameters using 14 years of real-world monitoring data from the Gorleben site.
The results demonstrate the efficiency and accuracy of the proposed approach. First, a sensitivity analysis with $16,384$ surrogate model calls was performed within $270$ seconds to identify four relevant material parameters. Afterwards, the GP-based surrogate model was calibrated based on convergence measurements, including the convergence of the global optimization in $29$ iterations, $1805$ model evaluations for gradient construction, and a duration of less than $1$ second.
This approach reduces the computational burden associated with traditional high-fidelity models and enables rapid, iterative updates to model parameters as new monitoring data becomes available. By enhancing the scalability and adaptability of numerical models, this work lays the foundation for integrating advanced surrogate modeling techniques into the management of deep geological repositories.
While the workflow demonstrated high efficiency and accuracy for the presented mechanical model, further developments are required to extend its applicability. Specifically, first, the GP-based surrogate model calibration and optimization method must be extended to account for higher-dimensional monitoring data, which could include extensometer and permeability measurements. Second, an alternative formulation would be developing a time-dependent surrogate model, which enables forecasting capabilities together with uncertainty propagation. Third, efforts have to be undertaken to account for the multi-physics nature of deep geological disposal. Apart from the purely mechanical model investigated here, a variety of models has been developed in the past that describe, e.g., transport of radio-nuclides through fluid flow, heat generation from high-level radioactive waste, or hydration of a sealing structure.
@@ Appendix A Appendix - Preliminary parameter selection
A preliminary parameter selection was performed in order to identify relevant strain components of TUBSsalt given the considered monitoring data. For this purpose, several damage-associated strain components were deactivated, in particular tertiary creep $\dot{\varepsilon}_{t}$, creep and shear failure $\dot{\varepsilon}_{n}$ and tension failure $\dot{\varepsilon}_{z}$. A comparison between $274$ gridpoint displacements at $40$ time instances is given in Figure A1 below, in which the vertical (Figure A1a) and horizontal (Figure A1b) displacements were simulated once with and once without damage strains. From the comparison, it can be observed that no deviation from the diagonal line is visible, which is also confirmed by the $R^{2}$ value close to $1.0$. Thus, the considered test case is dominated by creep mechanisms and softening, and post-failure strains are considered negligible.
@@ Appendix B Appendix - Material parameter
Table B1 contains parameter values and ranges for the constitutive model TUBSsalt. As reference serves a parameter set determined for Gorleben salt in [50]. However, it should be noted that the parameter set has been identified using data from various salt formations of Gorleben, and not only from the homogeneous z2HS2 area. As no parameters for healing are given in [50], $\eta_{v}$ and $m_{v}$ are obtained from [49].
Table B1: TUBSsalt material parameter under consideration of reference values from [50]. The parameters highlighted in red are directly read from experimental data. For the parameters highlighted in blue, a range of values is indicated in order to perform the sensitivity analysis.
Symbol Description Value or range Unit